OpenStudy (anonymous):

Which of the following is the equation of a line that passes through the origin? a. y=4/5x+8 b. y=x-1/5 c. y=4/3x d. y=-7x+3

OpenStudy (anonymous):

the answer is c) because the origin is (0,0) and all you have to do is sub in x=0 and y=0 into each of them and identify if the equation is true or not. y=4/3x 0 = 4/3(0) 0=0 so it is correct.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

no problem, medal thanks :D

OpenStudy (anonymous):

lol i had it right, just checking :) thnx though ahah

OpenStudy (anonymous):

haha kk.

OpenStudy (anonymous):

you can also double check by plotting it on your calculator.

OpenStudy (anonymous):

that's after you believe you are correct and have attempted doing the question, you shouldn't rely on the calculator too much.

OpenStudy (anonymous):

lol kk